Aubrie999:

So im so confused about math. So if you are solving a question like 15+c=2c-5 what do you do. Like how do you solve?

Hannahm:

try to skip it

Hannahm:

on the test

Aubrie999:

@hannahm wrote:
on the test
its not on the test it is an example

WanderingSamurai:

do you remember pemdas

WanderingSamurai:

:|

PoseidonGodOfTheSea:

2 + 2 = 4 see? Im good at meth LMAO jk

Olive2006:

Step 1: Simplify both sides of the equation. 15+c=2c−5 15+c=2c+−5 c+15=2c−5 Step 2: Subtract 2c from both sides. c+15−2c=2c−5−2c −c+15=−5 Step 3: Subtract 15 from both sides. −c+15−15=−5−15 −c=−20 Step 4: Divide both sides by -1. -c/-1 = -20/-1 c=20

Aubrie999:

@olive2006 wrote:
Step 1: Simplify both sides of the equation. 15+c=2c−5 15+c=2c+−5 c+15=2c−5 Step 2: Subtract 2c from both sides. c+15−2c=2c−5−2c −c+15=−5 Step 3: Subtract 15 from both sides. −c+15−15=−5−15 −c=−20 Step 4: Divide both sides by -1. -c/-1 = -20/-1 c=20
Thank you this really helps!

Olive2006:

@aubrie999 wrote:
@olive2006 wrote:
Step 1: Simplify both sides of the equation. 15+c=2c−5 15+c=2c+−5 c+15=2c−5 Step 2: Subtract 2c from both sides. c+15−2c=2c−5−2c −c+15=−5 Step 3: Subtract 15 from both sides. −c+15−15=−5−15 −c=−20 Step 4: Divide both sides by -1. -c/-1 = -20/-1 c=20
Thank you this really helps!
no problem, glad i helped

BarbieJoo:

So both sides have to equal the same thing right?? See even I don't understand, I have the same troubles